In my despair I cry to Thee (d–phrygian)

M: Martin Luther 1524 after Christ • T: Martin Luther 1524 after Christ due to Psalm 130 • E: N. Harnoncourt, G. Leonhardt, TCI 1994: Johann Sebastian Bach, Sacred Cantatas, BWV 38 •

Choral composition: February 26^{th}, 2010 after Christ: Norbert Südland

[figure: 1.; {In \atop In-}; {my \atop cline}; des- Thine; {pair \atop ear}; {I \atop un-}; {cry \atop to}; {to \atop my}; Thee, voice,; {Lord \atop and}; {God, \atop hear}; {hear \atop my}; Thou supp-; {me \atop li-}; {cal- \atop ca-}]

[figure: 1.; ling.; 2.; tions.; }; If; Thou; dost; mark; in-; i-; qui-; ties,; re-; mem-; ber]

[figure: all; our; fra-; il-; ties,; who,; Lord,; can; stand; be-; fore; Thee?]

In my despair I cry to Thee (e–phrygian)

M: Martin Luther 1524 after Christ • T: Martin Luther 1524 after Christ due to Psalm 130 • E: N. Harnoncourt, G. Leonhardt, TCI 1994: Johann Sebastian Bach, Sacred Cantatas, BWV 38 •

Choral composition: February 26^{th}, 2010 after Christ: Norbert Südland

[figure: 1.; {In \atop In-}; {my \atop cline}; des- Thine; {pair \atop ear}; {I \atop un-}; {cry \atop to}; {to \atop my}; Thee, voice,; {Lord \atop and}; {God, \atop hear}; {hear \atop my}; Thou supp-; {me \atop li-}; {cal- \atop ca-}]

[figure: 1.; ling.; 2.; tions.; }; If; Thou; dost; mark; in-; i-; qui-; ties,; re-; mem-; ber]

[figure: all; our; fra-; il-; ties,; who,; Lord,; can; stand; be-; fore; Thee?]